Utilizing the data shown in the contingency table 1 below, determine whether the relationship between types...

  1. Utilizing the data shown in the contingency table 1 below, determine whether the relationship between types of residential area and gun ownership is statistically significant. Once you have arrived at an answer for each question, please write a sentence or two interpreting the results, you may want to round the decimals to the nearest whole number (70 pts.).

  1. What are the alternative and null hypotheses (5pts)?
  1. Calculate the total column percentages (10pts.)

Gun Ownership

Types of Residential Area

Rural

Town

City

Total

Own Gun

218

206

131

555

(         %)

No Gun

375

379

235

989

(           %)

Total

593

585

366

1544

(           %)

  1. Calculate the expected frequency for each cell and explain the meaning of expected frequencies (15 pts).

Gun Ownership

Types of Residential Area

Rural

Town

City

Total

Own Gun

555

No Gun

989

Total

593

585

366

1544

  1. Calculate chi-square (χ 2) for each cell using to the formula (Oi − Ei )2/ Ei (15pt).

Gun Ownership

Types of Residential Area

Rural

Town

City

Total

Own Gun

No Gun

Total

    1. What is the Chi-square statistic(5 pt)? And what is the critical value for this case at 5% significance level, what information did you use to find it (10pts)?
    1. Draw a conclusion about your null hypothesis at 5% significance level, and explain how did you arrive that conclusion (10pt)
Answer #1

SOLUTION A) NULL HYPOTHESIS H0​: The two variables are independent

ALTERNATIVE HYPOTHESIS Ha​: The two variables are dependent

SOLUTION C) Expected Values RURAL TOWN CITY Total OWN GUN 593 x 555 = 213.157 585x 555 = 210.282 366x555 = 131.561 555 1544 1544 366989 =SOLUTION D)

Squared Distances RURAL TOWN CITY OWN GUN (218-213.157) 213.157 - 0.11 (206-210.282) 210.282 0.087 (131-131.561) 131.561 0.00SOLUTION E: The Chi-Squared statistic is computed as follows:

v_Y(i-bi) = 0.11+0.062 +0.087 +0.049 +0.002 +0.001=0.312

Based on the information provided, the significance level is α=0.05 , the number of degrees of freedom is df=(2−1)×(3−1)=2, Critical value of chi square is χ2= 5.991.

f) Decision about the null hypothesis: Since it is observed that χ2=0.312≤χc2​=5.991, it is then concluded that the null hypothesis is not rejected.It is concluded that the null hypothesis Ho is not rejected. Therefore, there is NOT enough evidence to claim that the two variables are dependent, at the 0.05 significance level.